# Segment as You Wish: Free-Form Language-Based Segmentation for Medical Images

### v1

@@ 4.4 Ablation Study on the Model Architecture
– Data Augmentation 0.883$\pm$0.012 0.930$\pm$0.017 0.289$\pm$0.011 0.328$\pm$0.019
– Trainable ImgEncoder 0.748$\pm$0.009 0.845$\pm$0.016 0.301$\pm$0.009 0.283$\pm$0.017
@@
We conducted an ablation study of FLanS on the FLARE22 dataset [39] to understand the contribution of each component, as presented in Table 4. Using an 80%-10%-10% train-validation-test split on the public FLARE22 training set, we evaluate the models’ performance on both the held-out test set and a transformed test set, which contained samples applied with random transformations from $O(2)$. Table 4 shows the prediction performance of FLanS and its variants, with components progressively removed. The results highlight that each component plays a crucial role in the model’s overall performance. Notably, while data augmentation improved the model’s robustness to random transformations, it slightly reduced performance on the canonical test set, as the model had to handle various transformations. However, by canonicalization network, the segmentation backbone focuses specifically on canonicalized medical images, thus achieving the best performance on both test sets.
4.5 Effective Understanding of Free-Form Text Prompts
@@
Figure 7: Left: t-SNE visualization of the free-form text prompt embedding space. Our method can effectively distiguish between different organ related queries. Right: Canonicalized CT scans from $D_{4}$ and $D_{8}$ canonicalization networks for a batch of randomly transformed scans from the FLARE22 dataset. Medical images can be successfully transformed back to an aligned canonicalization space.
4.6 Effectiveness of the Canonicalization
@@
The right side of Fig. 7 shows the canonicalized CT scans from $D_{4}$ and $D_{8}$ canonicalization networks for a batch of original scans from the FLARE22 dataset applied with random transformations from $O(2)$ group. As the group order of the canonicalization network increases, the scans become more consistently aligned to a particular canonical orientation. The canonicalization networks use a shallow architecture with three layers, a hidden dimension of 8, and a kernel size of 9, demonstrating that even a simple network with a larger kernel can effectively achieve canonicalization.
More importantly, applying canonicalization before feeding the scans into the main segmentation network and making the entire architecture equivariant or invarianthelps prevent confusion caused by positional terms in the organ name. A text-prompt segmentation model understands positional cues such as “left” vs “right” but it may get confused between the anatomical position and the organ’s appearance in the scan. For example, Fig. 6 shows segmentation predictions from models with and without canonicalization, given the anatomy-informed text prompt, “Highlight the right renal organ.” Since the CT scan is not in the standard orientation, the right kidney appears on the left side of the image. Without canonicalization, the non-equivariant model incorrectly segments the left kidney, which appears on the right side. Our model can make consistent predictions of the right kidney regardless of the scan’s orientation, allowing it to focus on learning the critical features of the organs.
5 Conclusion
In this work, we presented FLanS, a novel medical image segmentation model capable of handling diverse free-form text prompts, including both anatomy-informed and anatomy-agnostic descriptions. By integrating equivariance, our model ensures accurate and consistent segmentation across varying scan orientations, addressing a critical challenge in medical imaging. We also developed a RAG query generator for both realistic and synthetic prompt generation, and trained FLanS on over 100k medical images from 7 public datasets, covering 24 organ categories. FLanS outperforms baselines in both in-domain and out-of-domain tests, demonstrating superior language understanding and segmentation accuracy. Future works including extend FLanS to multi-organ segmentation tasks and further enhance RAG generator with multimodal data.
@@ Appendix A Equivariance and Symmetry
$f(\rho_{\text{in}}(g)(x))=\rho_{\text{out}}(g)f(x)$ (4)
for all $x\in X$ and $g\in G$, where $\rho_{\text{in}}\colon G\to\text{GL}(X)$ and $\rho_{\text{out}}\colon G\to\text{GL}(Y)$ are input and output representations [4]. Invariance is a special case of equivariance where the output does not change under the group action. This occurs when the output representation $\rho_{\text{out}}(g)$ is trivial. Figure 8 visualize how the equivariant and invariant networks work.
Figure 8: An equivariant model (left) ensures that its output transforms in a specific, predictable way under a group of transformations applied to the input, preserving the structure of the transformation (e.g., rotating the input results in a correspondingly rotated output). In contrast, an invariant model (right) produces an output that remains unchanged regardless of any transformations applied to the input from the same group.
@@
One of the primary approaches to incorporating symmetry into neural networks is through weight sharing [47, 9, 56]. This approach enforces equivariance by constraining the network’s architecture so that the weights are shared across different group elements. For example, in $G$-convolutions [11], the same set of weights is shared across the transformed versions of the input, ensuring that the network’s predictions remain consistent under those transformations. In a layer of $G$-steerable CNNs [60], a set of equivariant kernel bases is precomputed based on the input and output representations, and the convolution kernel used is a linear combination of this equivariant kernel basis set, where the coefficients are trainable. Similar approaches can also be used to develop equivariant graph neural networks [18]. These architectures directly modify the network’s layers to be equivariant, ensuring that each layer processes symmetries in a way that is aligned with the desired group. While powerful, this approach imposes architectural constraints, which may limit the flexibility of the network and prevent leveraging large pretrained models.
@@
An alternative to weight sharing is incorporating symmetry through canonicalization [25, 41], where, instead of modifying the network’s architecture to handle symmetries, the input data is transformed into a canonical form. In this approach, a separate canonicalization network, which is itself equivariant, preprocesses the input, transforming it into a standard, or canonical, representation. This canonicalized input is then passed to a standard prediction network that does not need to be aware of the symmetries. If the corresponding inverse transformation is applied to the output of the prediction network, the entire model becomes equivariant; otherwise, the model remains invariant. This method has several advantages. First, it does not require altering the architecture of the prediction network, allowing for the use of large pre-trained models without modification. Second, by ensuring that the input data is in a canonical form, the prediction network only needs to learn the mapping from the canonical input to the output, without needing to learn all transformed samples. This can lead to improved performance and robustness, especially in scenarios where the prediction task does not naturally align with the symmetry group or where architectural constraints might hinder performance. Thus, in our work, we leverage canonicalization to achieve equivariance in the segmentation task. By transforming the input into a canonical form using a simple equivariant canonicalization network, we ensure that our prediction network remains unconstrained and can fully utilize its capacity for learning without the need for architectural modifications. This approach offers the benefits of symmetry-aware processing while maintaining the flexibility and power of unconstrained neural network architectures.
@@ Appendix B Detailed Dataset Description
Table 5: Overview of the datasets used in this study.
@@
Figure 9: Distribution of labeled organs across the collected datasets. The image count for each organ and its corresponding ratio is marked in the plot.
@@
The overall structure of the test dataset is shown in Figure 10. It consists of 25% expert queries, 25% normal queries, and half synthetic queries. In total, we have 2880 (24 organs x10 queries x3 x2x2) text queries. Each of the queries is labeled with the correct organ name to segment. This will be used to evaluate the ability of our learned TextEncoder model to understand correct intentions based on free-form language description.
At the same time, the organ names are connected to another segmentation test set, which contains several (how many) medical images such as CT scans, MRIs, etc. And stand on the results of interest-category identification, we conduct further segmentation result analysis, including the normal segmentation precision study, and also the equivariant identified segmentation study.
@@
Figure 10: The Language Test Set for Verifying the Query Understanding Ability. It contains three aspects of components, real data - expert group, real data - non-expert group, and synthetic data.
Figure 11: Positional prompt dataset provider split, we take the slices with more than $\alpha$ labels, where we set $\alpha=8$ in this illustration (while 13 is the total label amount) as a split threshold, ensure that the slice used for training the label-agnostic provides sufficient semantics in the image content, such as left, upmost or largest, etc. Similarly, we process the other datasets such as BTCV and WORD.

### v2

@@ 4.4. Ablation Study on the Model Architecture
– Data Augm 0.883$\pm$0.012 0.930$\pm$0.017 0.289$\pm$0.011 0.328$\pm$0.019
– Trainable ImgEnc 0.748$\pm$0.009 0.845$\pm$0.016 0.301$\pm$0.009 0.283$\pm$0.017
@@
We conducted an ablation study of FLanS on the FLARE22 dataset (Ma et al., 2023) to understand the contribution of each component, as presented in Table 4. Using an 80%-10%-10% train-validation-test split on the public FLARE22 training set, we evaluate the models’ performance on both the held-out test set and a transformed test set, which contained samples applied with random transformations from $O(2)$. Table 4 shows the prediction performance of FLanS and its variants, with components progressively removed. The results highlight that each component plays a crucial role in the model’s overall performance. Notably, while data augmentation improved the model’s robustness to random transformations, it slightly reduced performance on the canonical test set, as the model had to handle various transformations. However, by canonicalization network, the segmentation backbone focuses specifically on canonicalized medical images, thus achieving the best performance on both test sets.
4.5. Understanding of Free-Form Text Prompts
@@
4.6. Effectiveness of the Canonicalization
Figure 8. The model without canonicalization incorrectly highlights the left kidney due to confusion between anatomical position (“right kidney”) and the organ’s appearance on the right side of the image.
Figure 9. The explanation behind the segmentation logic of two methods shown in the figure 8. FLanS trained with canonicalization module, is able to distinguish that, the question asking about ‘right renal’ is the actual ‘right’ to the human body, not the observation side, since ‘right kidney’ is essentially different from ‘right side kidney’, and our model handle it by a well-learned canonicalization space and transforming the input image $x$ to the $x^{\prime}$ based on the body standard orientation. The segmentation without such a canonicalization module can only directly process the input image, thus making mistakes when it comes to language ambiguity.
@@ 4.6. Effectiveness of the Canonicalization
The right side of Fig. 7 shows the canonicalized CT scans from $D_{4}$ and $D_{8}$ canonicalization networks for a batch of original scans from the FLARE22 dataset applied with random transformations from $O(2)$ group. Increasing the group order of the canonicalization network improves alignment to a canonical orientation. Using a simple three-layer network with a hidden dimension of 8 and a kernel size of 9 effectively achieves this alignment.
Discussion: Why not just perform augmentation but apply canonicalization? Canonicalization ensures equivariance, resolving positional confusion that simple augmentation cannot. For instance, in Fig. 8, a model using only augmentation misidentifies the left kidney as the right one when given the prompt “Highlight the right renal organ,” because the input scan is not in the standard orientation and the right side kidney appears on the left side of the image. Canonicalization addresses this by aligning scans to a consistent orientation, enabling accurate segmentation regardless of the scan’s initial positioning. Specifically, FLanS trained with canonicalization module, is able to distinguish that, the question asking about ‘right renal’ is the actual ‘right’ to the human body, not the observation side, since ‘right kidney’ is essentially different from ‘right side kidney’, and our model handle it by a well-learned canonicalization space and transforming the input image to the canonicalized one based on the body standard orientation and then perform the segmentation as shown in Figure. 9.
5. Conclusion
We presented FLanS, a medical image segmentation model effectively handling diverse free-form prompts. By dividing the real-world query into anatomy-informed and anatomy-agnostic scenarios, we could successfully train a flexible foundation model by proposing a RAG-based generator for description generation and the equivariance module for efficient image feature capturing. Through this design, FLanS learns a strong understanding of the free-form language descriptions that promote arbitrary segmentation across varying orientations with high accuracy. The model has the potential to generalize to more modalities and organ classes for real-world clinical deployment.
6. Limitations and Future Work
This paper explores a solution that bridges the language flexibility and difficult segmentation tasks in the medical image domain. Although we demonstrate that the current model performs well in CT scans with various text prompts, we found that its performance in other modalities (e.g., MRI or Ultrasound) is still limited. Besides, the current model’s ability in processing positional hints is fixed to types (latest, smallest, or left-most, etc), to advance the flexibility, possible explorations can be made on combinational prompts, e.g., Size + Positional is an interesting direction.
Acknowledgments
The work was partially supported by NSF awards #2421839. The views and conclusions in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Equivariance and Symmetry
(5) $f(\rho_{\text{in}}(g)(x))=\rho_{\text{out}}(g)f(x)$
for all $x\in X$ and $g\in G$, where $\rho_{\text{in}}\colon G\to\text{GL}(X)$ and $\rho_{\text{out}}\colon G\to\text{GL}(Y)$ are input and output representations (Bronstein et al., 2021). Invariance is a special case of equivariance where the output does not change under the group action. This occurs when the output representation $\rho_{\text{out}}(g)$ is trivial. Figure 10 visualize how the equivariant and invariant networks work.
Figure 10. An equivariant model (left) ensures that its output transforms in a specific, predictable way under a group of transformations applied to the input, preserving the structure of the transformation (e.g., rotating the input results in a correspondingly rotated output). In contrast, an invariant model (right) produces an output that remains unchanged regardless of any transformations applied to the input from the same group.
@@
One of the primary approaches to incorporating symmetry into neural networks is through weight sharing (Satorras et al., 2021; Cohen et al., 2018; Wang et al., [n. d.]). This approach enforces equivariance by constraining the network’s architecture so that the weights are shared across different group elements. For example, in $G$-convolutions (Cohen and Welling, 2016), the same set of weights is shared across the transformed versions of the input, ensuring that the network’s predictions remain consistent under those transformations. In a layer of $G$-steerable CNNs (Weiler and Cesa, 2019), a set of equivariant kernel bases is precomputed based on the input and output representations, and the convolution kernel used is a linear combination of this equivariant kernel basis set, where the coefficients are trainable. Similar approaches can also be used to develop equivariant graph neural networks (Geiger and Smidt, 2022). These architectures directly modify the network’s layers to be equivariant, ensuring that each layer processes symmetries in a way that is aligned with the desired group. While powerful, this approach imposes architectural constraints, which may limit the flexibility of the network and prevent leveraging large pretrained models.
@@
An alternative to weight sharing is incorporating symmetry through canonicalization (Kaba et al., 2022; Mondal et al., 2023), where, instead of modifying the network’s architecture to handle symmetries, the input data is transformed into a canonical form. In this approach, a separate canonicalization network, which is itself equivariant, preprocesses the input, transforming it into a standard, or canonical, representation. This canonicalized input is then passed to a standard prediction network that does not need to be aware of the symmetries. If the corresponding inverse transformation is applied to the output of the prediction network, the entire model becomes equivariant; otherwise, the model remains invariant. This method has several advantages. First, it does not require altering the architecture of the prediction network, allowing for the use of large pre-trained models without modification. Second, by ensuring that the input data is in a canonical form, the prediction network only needs to learn the mapping from the canonical input to the output, without needing to learn all transformed samples. This can lead to improved performance and robustness, especially in scenarios where the prediction task does not naturally align with the symmetry group or where architectural constraints might hinder performance. Thus, in our work, we leverage canonicalization to achieve equivariance in the segmentation task. By transforming the input into a canonical form using a simple equivariant canonicalization network, we ensure that our prediction network remains unconstrained and can fully utilize its capacity for learning without the need for architectural modifications. This approach offers the benefits of symmetry-aware processing while maintaining the flexibility and power of unconstrained neural network architectures.
@@ Appendix B Detailed Dataset Description
Table 5. Overview of the datasets used in this study.
@@
Figure 11. Distribution of labeled organs across the collected datasets. The image count for each organ and its corresponding ratio is marked in the plot.
@@
The overall structure of the test dataset is shown in Figure 12. It consists of 25% expert queries, 25% normal queries, and half synthetic queries. In total, we have 2880 text query entries from 24 organs. Each of the queries is labeled with the correct organ name to segment. This will be used to evaluate the ability of our learned TextEncoder model to understand correct intentions based on free-form language description. At the same time, the organ names are connected to another segmentation test set, which contains various medical images such as CT scans, MRIs, etc. And stand on the results of interest-category identification, we conduct further segmentation result analysis, including the normal segmentation precision study, and also the equivariant identified segmentation study.
Appendix C Training Details
In the training process, we provide details of the configuration files and instructions below:
All experiments were conducted on an AWS ml.p3dn.24xlarge instance equipped with 8 V100 GPUs, each with 32 GB of memory. We used a batch size of 16 and applied the CosineAnnealingLR learning rate scheduler, initializing the learning rate for all modules at 0.0001. The AdamW optimizer was employed for training. A small $D_{8}$ equivariant canonicalization network was used, consisting of 3 layers, a hidden dimension of 8, and a kernel size of 9. To maintain consistency across the input and output formats, all scans from different datasets were resized to 1024$\times$1024 and both predicted and ground truth masks were resized to 256$\times$256 for fair comparison. For images with a single channel, the channel was duplicated to 3. All models’ performance on the test sets is reported using both the Dice coefficient (Taha and Hanbury, 2015) and normalized surface distance following the existing work (Heimann and Meinzer, 2009).
Appendix D Attention Map
In this section, we demonstrate the attention map to visualize the correlation between text embedding and image embedding.
These attention maps based on autonomy-agnostic prompts not only address the first reviewer’s query about plotting the relationship between text and images but also counter the second reviewer’s concern that ’our method seems capable of inferring which organ is desired given a text prompt and images.
@@ Appendix D Attention Map
Figure 12. The Language Test Set for Verifying the Query Understanding Ability. It contains three aspects of components, real data - expert group, real data - non-expert group, and synthetic data.
For example, consider the attention map generated with the prompt ‘segment the part located at the topmost portion’ it does not highlight just one organ. Instead, all organs at the top are highlighted. This demonstrates that our model is not merely overfitting data to infer a specific organ; rather, it has a deep understanding of the text and its relationship to the medical image.
Figure 13. The examples of attention maps in FLanS segmentation tasks. The attention maps are computed based on the scaled product of text embedding from the text encoder and image embedding from the image encoder.
Appendix E More Use Case Demonstrations
In this section, we provide more examples of segmentation results, ranging from the Anatomy-Informed to Anatomy-Agnostic ones. In Anatomy-Informed segmentation, we conduct two versions of illustration, first, we show the simplest (organ name is explicitly described in the prompt) segmentation, as shown in Fig. 16, and to add on more complexity and showcase how FLanS is beneficial for real-world clinical use cases, we take the diagnosis data from pseudonymized real EMR (electronic medical record) data in Fig. 14, FLanS is able to detect the organ from the long and redundant descriptive symptoms and provide accurate segmentation, this is especially useful for providing diagnosis assistance based on doctors’ notes. And last, we show size late Anatomy-Agnostic segmentations as in Fig. 15.
Figure 14. The implicit Anatomy-Informed demonstration on a deployed version of the FLanS model, it provides real-time inference ability and can be robust to any format type of the prompts, either lengthy or redundant, it can still perform effectively to identify the most cared organ. In this image, two examples of real EMR record data are provided (personal information such as name and age is pseudonymized.) We could observe that, even though sometimes the actual label is not explicitly described as in the upper half of this image, the model can perform segmentation accurately because it has seen the semantic similar corpus in the training period and aligned these symptom-related texts with the correct segmentation area.
Figure 15. The Anatomy-Agnostic demonstration on a deployed version of the FLanS model, this feature is specially designed for a larger group of users who care about the medical image scannings, but lack the professional background knowledge, such as students, and patients. This supports the segmentation with the size relevant or positional relevant prompts, the above figures show that FLanS is able to successfully segment those largest, sand mallest organs in the provided scans.
Figure 16. The explicit Anatomy-Informed demonstration on a deployed version of the FLanS model, in this figure, we showcase four examples of segmentation results by organ names relevant prompts. In the prompt content, it mentions the organ names, so as to instruct the segment action. As in the ① and ②, the model segment spleen and aorta successfully. And as shown in the ③ and ④, the model is also able to distinguish the right kidney and left kidney regardless of the angle and position of the scan is taken, enforced by the canonicalization module as introduced in the Section 3.3.
Figure 17. This image provides an extra analysis on the corner case detection. As shown in this image, there exist situations in blurry, low-quality metadata leading to the invisibility of an organ, or that an organ does not exist. FLanS is able to detect such cases and provide feedback that The organ queried is likely not existing in this image!. This is realized by a filter layer of function upon the predicted Probability for an organ area, the threshold is set to $\alpha=0.5$. One can easily tune the parameter based on the actual required confidence of the FLanS.